Find the value of y it's supplement is 4times it's compliment

Answer by mananth(16946)

Find the value of y it's supplement is 4times it's compliment

supplement of y = (180-y)
complement of y = (90-y)
supplement of y = 4 times complement of y
180-y=4(90-y)
180-y=360-4y
3y=360-180
3y=180
y=60